OpenStudy (anonymous):

French Co. marks up its goods 45% on cost. What is French's equivalent markup on selling price? Round to nearest hundredth percent

OpenStudy (anonymous):

can anyone help me?

OpenStudy (kropot72):

Let CP = cost price Let SP = Selling price = CP + 0.45CP Mark up = 0.45CP Ratio of Mark up to Selling price = 0.45CP/(CP + 0.45CP) = 0.45/(1+0.45)

OpenStudy (kropot72):

Can you follow the working and find the answer?

OpenStudy (anonymous):

ugh im kinds confused. what wud the cp be than?

OpenStudy (kropot72):

CP can be any price. Notice that the CP cancels out in the working.

OpenStudy (anonymous):

so would it be 1.65?

OpenStudy (kropot72):

No. The ratio of Mark up to Selling price is 0.45/(1+0.45)=0.45/1.45=0.3103. So what is the percentage mark up on the selling price?

OpenStudy (anonymous):

i have no clue >.<

OpenStudy (anonymous):

im sorry. its 1 am. i have 2 questions left and im really confused and not good at math

OpenStudy (kropot72):

It is 0.3103 multiplied by 100 = 31.03%

OpenStudy (anonymous):

so since it says roud to nearest hundreth percent it wud be 31% or wud it just be 31.03%

OpenStudy (kropot72):

One hundreth of a percent is .01%. So 31.03% is correct.

OpenStudy (anonymous):

oooooh okay thank you sooo much